OpenStudy (anonymous):

h(x)=sqrt4-x find h(0) + h(-5) what is h???

OpenStudy (agent0smith):

h is the function, h(x). Means h is a function of x. Think of h(x) as being like y, like this \[\large y = h(x) = \sqrt {4-x}\] h( ) <==== whatever is inside the brackets, plug that in in place of x in sqrt(4-x). h(0) means... what is h equal to, when x is 0? or think of it as what is y equal to, when x is 0.

OpenStudy (anonymous):

y is equal to 5?

OpenStudy (anonymous):

oh its -1 so its undefined

OpenStudy (anonymous):

?

OpenStudy (agent0smith):

wait, what? no, y is defined for both of these x values h(0) + h(-5) ^this means plug in x=0 ^ this means plug in x=-5 h(0) = sqrt (4-0) h(-5) = sqrt (4- (-5))

OpenStudy (anonymous):

would my answer be undefined ?

OpenStudy (agent0smith):

No... simplify everything inside the sqrt before taking the square root.

OpenStudy (agent0smith):

sqrt (4-0) = ??? sqrt (4- (-5)) = ??? then add the results together

OpenStudy (anonymous):

my answer is 5 - i completely forgot to do the sqrt of 4 i hate when I miss simple things like that!

OpenStudy (agent0smith):

Good job :)

OpenStudy (anonymous):

=) thank you